Determine the height of the flag at the top of the flagpole. Show your work. 5m 3m
Answer #1

Let the height of the flag is x

Consider flagh as a Right triangle

Using pythegorean theoram.

x=\sqrt{5^2-3^2}=\sqrt{25-9}=\sqrt{16}=4

Height of flag is 4 m